OpenStudy (anonymous):

cot (3pie/8). How do I plug in this problem into formula: tan(A/2) = √1-cosA/1+cosA ??

OpenStudy (anonymous):

\[\large \cot(\frac{3\pi}{8})=\frac{1}{\tan(\frac{3\pi}{8})}\] \[\large =\frac{1}{\tan(\frac{1}{4}\times \frac{3\pi}{2})}\]

OpenStudy (anonymous):

how did you get tan (1/4 x 3pie/2) ???

OpenStudy (anonymous):

? Please take a minute to read that I separated the denominator.